OpenStudy (anonymous):

Solve √3x+10 =x for x.

OpenStudy (e.mccormick):

OK. You need to get all the Xes together in solving for x. If a square shows up, which may happen since there is a root, that can result in more than one answer.

OpenStudy (e.mccormick):

What did you try to start and how far did you get?

OpenStudy (anonymous):

I couldn't try because I didn't know how to start

OpenStudy (anonymous):

\[x=10/(1-\sqrt{3})\]

OpenStudy (e.mccormick):

AH. OK! Now, is that square root over the whole thing? All I see is it next ot the 3. So does it get the 10 also?

OpenStudy (anonymous):

Yes, except for the =x part

OpenStudy (e.mccormick):

OK. Then the fist step is to square both sides of the equation. This will let you get rid of the root.

OpenStudy (anonymous):

that is what i think @e.mccormick it is next to the 3 except if not then the answer is different

OpenStudy (anonymous):

3x+10=x^2

OpenStudy (anonymous):

i took the x values to one side

OpenStudy (anonymous):

@ballerina789 but if it is \[\sqrt{3x}\] then you have to square both sides

OpenStudy (anonymous):

if it is \[\sqrt{3}x\] then all you have to do is take all the x components to one side

OpenStudy (anonymous):

the √ goes over the whole equation. So when I square it it would equal: 3x+10=x^2

OpenStudy (e.mccormick):

Yes.

OpenStudy (e.mccormick):

Do you know where to go with that?

OpenStudy (anonymous):

move everything to one side?

OpenStudy (e.mccormick):

Yep. And a positive \(x^2\) is nice. So I would move the 3x+10 side.

OpenStudy (anonymous):

x^2-3x+10=0

OpenStudy (e.mccormick):

Now, can you factor that?

OpenStudy (anonymous):

x^2-3x-x+10=0 ?

OpenStudy (e.mccormick):

Well, what you are looking for is something like:\[(x\pm\,?\,)(x\pm\,?\,)\]

OpenStudy (anonymous):

(x-5)(x+2)

OpenStudy (anonymous):

x=5 x=-2

OpenStudy (anonymous):

correct?

OpenStudy (e.mccormick):

Well, there is one other thing to do. You are almost there. Those arew correct, but you need to find x and verify it.

OpenStudy (anonymous):

ok?

OpenStudy (e.mccormick):

You see, because we started with \(\sqrt{3x+10}\) anything that makes the part under the root negative is NOT an answer.

OpenStudy (e.mccormick):

So make sure that nothing you did would result in that being negative.

OpenStudy (anonymous):

oh so the answer would only be 5

OpenStudy (e.mccormick):

On a graph, only 5 works. In this case, -2 would mathematically work. But it does not work on a graph. Let me show you why: https://www.desmos.com/calculator/a67kx3emql

OpenStudy (anonymous):

ahhhh, ok. So the answer to the problem would be x=5

OpenStudy (e.mccormick):

Yah. Now, the odd thing is caused by squaring it. Doing a square can add answers that do not work. When we squared we got: \(x^2-3x-10\). Let me see if I can add that to the graph.

OpenStudy (e.mccormick):

Connection issues.... was going to show you one last thing. It will take a min. But I want you to get the point.

OpenStudy (e.mccormick):

I added the \(x^2-3x-10\) to the graph. https://www.desmos.com/calculator/pvjui3pw5b Now, on the square, in orange, you are looking at the x intercepts. It shows two. However, the graph of the original equations is looking at the intersection. The true test to see if the answers are right or wrong is to actually test them and do the math. \(\sqrt{3(-2)+10}=-2\Rightarrow \sqrt{-6 +10}=-2 \Rightarrow 2=-2 \) which as you correctly saw, was false. In fact, anything under a root = to a negative is false, until you get to use complex numbers. So as long as the teacher has not said specifically, "Complex solutions," we assume real solutions and eliminate negatives like this. \(\sqrt{3(5)+10}=5\Rightarrow \sqrt{15 +10}=5 \Rightarrow 5=5 \) Now that tests as true. So that is a real solution and useable. The moral of the story is simple: Test answers for roots!

OpenStudy (e.mccormick):

https://www.desmos.com/calculator/pvjui3pw5b Now, on the square, in orange, you are looking at the x intercepts. It shows two. However, the graph of the original equations is looking at the intersection. The true test to see if the answers are right or wrong is to actually test them and do the math. \(\sqrt{3(-2)+10}=-2\Rightarrow \sqrt{-6 +10}=-2 \Rightarrow 2=-2 \) which as you correctly saw, was false.